Predetermined Factory Overhead Rate

Poehling Medical Center has a single operating room that is used by local physicians to perform surgical procedures. The cost of using the operating room is accumulated by each patient procedure and includes the direct materials costs (drugs and medical devices), physician surgical time, and operating room overhead. On January 1 of the current year, the annual operating room overhead is estimated to be:

Disposable supplies $281,800
Depreciation expense 50,800
Utilities 29,600
Nurse salaries 423,200
Technician wages 138,600
Total operating room overhead $924,000

The overhead costs will be assigned to procedures, based on the number of surgical room hours. Poehling Medical Center expects to use the operating room an average of eight hours per day, seven days per week. In addition, the operating room will be shut down two weeks per year for general repairs.

a. Determine the predetermined operating room overhead rate for the year.
$ per hour

b. Bill Harris had a 5-hour procedure on January 22. How much operating room overhead would be charged to his procedure, using the rate determined in part (a)?
$

c. During January, the operating room was used 194 hours. The actual overhead costs incurred for January were $62,600. Determine the overapplied operating overhead or underapplied operating overhead for the period. Enter your answer as a positive number.
$

Overapplied or Underapplied
Answer #1

a) Estimated hour = (52-2)*7*8 = 2800 Hour

Predetermine overhead rate = estimated overhead/estimated hour = 924000/2800 = 330 per hour

b) Operating overhead applied = 330*5 = $1650

c) Applied overhead = 194*330 = 64020

Actual overhead = 62600

Over applied overhead = 64020-62600 = 1420